# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v1

@@ B.2 Proof of Lemma 4
$\displaystyle\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (19)
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
@@
$\displaystyle\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (20)
$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
@@
$\displaystyle\quad\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (21)
$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
@@
$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
The inner product can produce a similar term. According to Lemma 3,
$\mathbb{E}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}[H_{m}({\mathbf{y}}^{(k)},\nabla f({\mathbf{x}}^{(k)},\ldots,\nabla f({\mathbf{x}}^{(k+m-1)}))].$
Therefore, we have
$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (22)
$\displaystyle=-\alpha\mathbb{E}\left[\left\langle\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),\nabla f(\hat{x}^{(k)})\right\rangle\right]$
$\displaystyle\quad+\mathbb{E}\left[\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right){\mathbf{y}}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$
$\displaystyle\leq\frac{1}{cmL}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}].$
Plug (20) and (22) into (19), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (23)
@@
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
For $2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
Take (24) into (23), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (25)
@@
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
We can further apply Jensen inequality on the last term in (25), i.e.
$\displaystyle\leq(\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2})^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
$\displaystyle\leq s_{B}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
Therefore, (25) can be further simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
By taking $\alpha\leq\min\{\frac{1}{8cmL},\frac{c}{8n^{2}s_{B}^{2}L}\}$, this can be further simplified to
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{L}{16m}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{L}{16m}\sum_{j=1}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
@@
B.3 Proof of Lemma 5
We consider how to measure the difference of ${\mathbf{x}}$. Use the fact that
$\displaystyle{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
We decompose ${\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ and apply the Cauchy-Schwarz inequality:
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]\leq 3\left(\sum_{i=1}^{m-1}\|A^{i}-A_{\infty}\|_{2}^{2}\right)\|\Delta_{x}^{(k)}\|_{F}^{2}$ (27)
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\|\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\|_{2}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}+3\alpha^{2}s_{B}^{2}L^{2}\Delta_{2}^{(k)},$
where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness.
For ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, similarly, we decompose it into
$\displaystyle{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})+\Delta_{x}^{(k)},$
which finally produces
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]\leq 4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (28)
$\displaystyle\quad+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4\alpha^{2}s_{B}^{2}L^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]+4m\|\Delta_{x}^{(k)}\|_{F}^{2}.$
Inequalities (27) and (28) tell us
$\displaystyle\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}$
where $v_{1}=3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$, $v_{2}=4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4m\|\Delta_{x}^{(k)}\|_{F}^{2}$ are positive numbers. To proceed on, we assume $\alpha\leq\frac{1}{6s_{B}L}$, so that the inversion of the $2\times 2$ coefficient matrix satisfies
$\displaystyle\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\leq{\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}}^{-1}\leq\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}.$
Therefore, we have
$\displaystyle{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{10m}\\ \frac{L}{10m}\end{pmatrix}}^{\top}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}.$
This gives that
$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2},$
which finishes the proof of Lemma 5.
B.4 Proof of Lemma 6
In this section, we provide some inequalities for the consensus error $\Delta_{x}$ and $\Delta_{y}$. These inequalities will only be used in the final part of proof of our main theorem.
Lemma 8 (Consensus of x).
Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, we have
$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]$
$\displaystyle\quad+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2}$

### v2

@@ B.1 Proof of Lemma 4
$\displaystyle\frac{\alpha cm}{2}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (25)
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
@@
$\displaystyle\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (26)
$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
@@
$\displaystyle\quad\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (27)
$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}_{k}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
@@
$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
The inner product can produce a similar term:
$\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}_{k}[H_{m,k}].$
Using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]=-\alpha\left\langle\mathbb{E}_{k}[H_{m,k}],\nabla f(\hat{x}^{(k)})\right\rangle$ (28)
$\displaystyle=-\alpha\left\langle\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))],\nabla f(\hat{x}^{(k)})\right\rangle$
$\displaystyle\quad+\alpha\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle$
[…]
$\displaystyle\quad+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\leq\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]+\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\quad+\frac{2\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
where the last inequality uses $\|\pi_{A}\|\leq 1$ and $\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}\leq s_{B}$. Plug (26) and (28) into (25), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (29)
$\displaystyle\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+(2\alpha^{2}Ls_{B}^{2}+\frac{2\alpha s_{B}^{2}}{cm})\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{2\alpha}{cm})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
For $c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
Take (30) into (29), and suppose $\alpha\leq\frac{1}{4cmL}$, we obtain that
@@
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
We can further apply Cauchy’s inequality on the last term in (31), i.e.
$\displaystyle\leq\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
$\displaystyle\leq s_{B}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (31) can be further simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
@@
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{j=1}^{k}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
@@
B.2 Proof of Lemma 5
In this section, we consider how to measure the difference of ${\mathbf{x}}$. To be more specific, we mainly consider the relationships between the following 44 44 terms:
$\Delta_{1}^{(k)},\Delta_{2}^{(k)},\mathbb{E}_{k}[\|\Delta_{x}^{(k)}\|_{F}^{2}],\Delta_{y}^{(k)},\quad k\in[m^{*}].$
Lemma 7 will give upper bounds of $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ using other terms, while Lemma
Lemma 7.
Suppose Proposition 2 and Assumption 34 hold. When $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold for any positive integers $k$ and $m$.
$\displaystyle\Delta_{1}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (32)
$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$
$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (33)
$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$